# Reconciling aging and slip state evolutions from laboratory-derived canons of rate-and-state frictionF

### v2

@@ 5.3 Potential detection of $\beta$ from an experiment
As far as $\dot{\theta}=f(V,\theta)$, we have theoretically shown that the aging-slip switch must occur around $\Omega=\beta$ to reproduce the three canonical behaviors confirmed by experiments so far. Such evolution laws involving an aging-slip switch are necessarily accompanied by some side effects violating the canons (§4.1; also, see §4.2.1). Among them, relatively accessible by experiments may be the behavior following a large negative velocity step ($V_{\rm after}/V_{\rm before}\ll\beta$), which is expected to exhibit too rapid strengthening compared to the prediction by the slip law, that is, the VS canon (e.g., Fig. 11a). Although existing VS tests generally obey the slip law as we presumed as the VS canon throughout the present development, a slight deviation has been reported in fact. Bhattacharya et al. (2015) realized that an interesting behavior of the Nagata law (Nagata et al. 2012), which is the aging law plus a stress-dependent term $-cd\tau$ representing an experimentally detected instantaneous effect of the applied shear stress to weaken the interface; that is, the Nagata law changes from the aging law to the slip law for large velocity steps. They further showed that some VS datasets involving multiple steps with various levels of $V_{\rm after}/V_{\rm before}$ (ranging $10^{-2}$ to $10^{2}$) are fit well by the Nagata law with $c=50$, contrasting that the slip law cannot fit the large negative VS. This finding of Bhattacharya et al. (2015) implies that the involvement of the aging-slip switch improves fitting to the experiments.
As Bhattacharya et al. (2015) themselves point out, assuming $c$ as large as 50 is unreasonable as the shear-stress effect of the Nagata law, given that Nagata et al. (2012) obtained $c=$2 by direct measurement of $d\Phi/d\tau$. However, from the viewpoint of the aging-slip switch, their Nagata law fits suggest the aging-slip switch according to $\Omega$ we deduced from the canons. Below, we try fitting the data shown in Bhattacharya et al. (2015) using the proposed modified composite law. Of course, making a formal statistical conclusion on the aging-slip switch according to $\Omega$ requires to demonstrate the increase in likelihoods sufficient to counter the increase in the number of parameters. The slight deviation found in the $V_{\rm after}/V_{\rm before}$ range of the data used in Bhattacharya et al. (2015) is probably not enough to conclude the existence of the additional parameter (super large $c$, or more reasonably, $\beta$) (Bhattacharya et al. 2015). Therefore, a definitive conclusion must be postponed until more suitable data becomes available. Further detailed discussion shall be shown at the end of this subsection, where we also consider Bhattacharya et al. (2022) involving exploration of a further lower range of $\Omega$.
@@
In the above, we have pointed out that the modified-composite law with $\beta=0.01$ explains the data of Bayart et al. (2006) for $V_{\rm after}/V_{\rm before}=0.01$, which Bhattacharya et al. (2015) have found to be basically slip-like but somewhat skewed toward aging-like behaviors, which are approximated by the Nagata law with excessively large $c$-values. At the same time, we admit that some datasets do not support the aging-slip switch. First, Bhattacharya et al. (2015) note that the above-mentioned better fit of the Nagata law than the slip law, which supports the aging-slip switch, was not evident in one of the three datasets they examined. Second, Bhattacharya et al. (2022) have performed VS tests, planned for capturing the composite-law-like aging-slip switch, strategically including quite low nm/s-order slip rates and small and large steps. Although they have conducted VS tests with $V_{\rm after}/V_{\rm before}\geq 10^{-3.5}$, they have not reported the aging-slip switch. It implies $\beta$ for their samples is below $10^{-3.5}$ if exits. Still, this observation of Bhattacharya et al. (2022) may originate from the long-term trend  (similar to Blanpied et al., 1998b) in their data which are not fully taken into their analysis. They actually reported that the observed slip-strengthening behaviors are even slower than the slip law. Roughly removing that long-term trend by a linear fit, we found that the modified composite law with $\beta=0.01$ well fits their $V_{\rm after}/V_{\rm before}=10^{-3.5}$ data, although convincing discussions about detrending are necessary in this regard. Thus, we deem $10^{-3.5}<\Omega<10^{-2}$ still under debate.
After all, we do not know much about the $\beta$ value, because VS tests with $V_{\rm after}/V_{\rm before}\leq 10^{-2}$ is still scarce. However, we emphasize that the existence of $\beta$ has been deduced from the reconciliation of the three canons under the premise of $f(V,\theta)$, and our SHS canon (significant log-t healing in low-$\tau_{\rm hold}$ SHS) is set at $\Omega$ far-far below steady states ($\Omega_{\rm SS}=1$), that is, $\Omega\to 0$, much lower than $10^{-3.5}$. As the proposers, we reckon that VS tests with further small $V_{\rm after}/V_{\rm before}$ will find the expected crossover. Also, note that the value of $\beta$, like the value of $A$ and $B$, can depend on various experimental conditions other than the variables of friction law ($\tau$, $V$, $\Phi$), including rock types, temperature, normal stress, and gouge conditions. Hence, $\beta$ must be considered a fitting parameter, which should be experimentally determined under each condition, as $A$ and $B$ are. For prudence, we still regard the aging-slip switch and its side effects as an unconfirmed theoretical prediction of the present work, and we would like to present the data analysis shown in Fig. 12 as cursory support to our proposal.
@@ 6 Conclusion
In the second step (§4.2.2), we have further shown that the aging-slip crossover occurring around $\Omega=\beta$ is an inevitable consequence of the three canons even when evolution laws are sought in a more general scope of $\dot{\theta}=f(V,\theta)=f_{t}+f_{\delta}V$ with the $f=\psi(\Omega)$ ansatz removed (eq. 86).
@@ Acknowledgements.
We would like to thank Dr. Hugo Perfettini, Dr. Paul Antony Selvadurai, Dr. So Ozawa, Dr. Bunichiro Shibazaki, Dr. Yukitoshi Fukahata, Dr. Takeshi Iinuma, and Dr. Takane Hori for valuable discussions and encouragement. We also acknowledge careful reviews and insightful comments from Dr. Allan Rubin, Dr. Shiqing Xu, Dr. Baoning Wu, and an anonymous reviewer. This research was supported in part by JSPS KAKENHI Grant Numbers 22KJ1658 and 23K19082.
@@ Appendix A Predictions of different evolution laws during quasi-stationary contact
This appendix presents the behaviors expected of different evolution laws discussed in the present paper (i.e., the aging, slip, composite, and modified composite laws) during constant-$\tau$ SHS tests explained below. In this type of SHS test, following a steady-state sliding at $V=V_{\rm prior}$, shear stress is abruptly reduced by $\Delta\tau$ ($>0$) to a designated level $\tau_{\rm hold}$ ($<\tau_{\rm SS}(V_{\rm prior})$) and held constant for a designated time duration $t_{\rm h}$. Regarding the log-t healing behavior, we focus on constant-$\tau$ SHS tests because the NSC can be ensured throughout the hold period by setting the $\tau_{\rm hold}$ sufficiently low.
For all the four evolution laws considered here, initial conditions at the beginning of a quasi-stationary hold period (defined as $t=+0$) are given as follows:
$V|_{\rm t=0+}$ $=:V_{\rm init}=V_{\rm prior}e^{-\Delta\tau/A}$ (A.1)
$\theta|_{\rm t=0+}$ $=:\theta_{\rm init}=L/V_{\rm prior}$ (A.2)
Equation (A.1) derives from eq. (1), irrespective of the evolution law here considered. Note that eq. (A.2) assumes each evolution law satisfies $\Omega_{\rm SS}=1$ (eq. 31). This is not always the case, for instance, for the composite law; associated remarks will appear in §A.1 and §A.4. Throughout Appendix A, we also assume $D=L=D_{\rm c}$ and $\theta\gg\theta_{\rm X}$.
@@
where $\Omega_{\rm init}:=\Omega|_{t=0+}$. Equation (A.4) shows that $V$ decreases as the interface strengthens under the constant $\tau_{\rm hold}$ imposed. Equation (A.5) shows that $\Omega$ remains constant if $A=B$, monotonically increases if $A>B$, and monotonically decreases if $A<B$. Using eqs. (A.4) and (A.5), we will analyze the evolution of $\theta$ (and $\Phi$ via eq. 11) for each evolution law in §A.2A.5, after cautioning a problem of constant-$\tau$ SHS tests for interfaces with $A>B$ in §A.1. We focus on behaviors under the NSC by setting $\tau_{\rm hold}$ sufficiently low such that
@@
Figure A.1 summarizes the exact evolutions obtained by numerical integration for all cases of the present concern. We will explain each case through analytical evaluations and discuss how it agrees or disagrees with the canonical behavior of log-t healing (§3.3 and 3.4).
@@ A.1 Healing saturation — a flaw of constant-$\tau_{\rm hold}$ tests for frictional interfaces with $A>B$
In general, $\Phi_{\rm SS}(V)$ increases during a hold period because $V$ decreases (eq. A.4) as healing proceeds (i.e., $\theta$ increases). When $A\leq B$, while both $\Phi$ and $\Phi_{\rm SS}$ increases with time, $\Phi$ never catches up with $\Phi_{\rm SS}$; in fact, eq. (A.5) guarantees that $\Phi$ more and more overwhelms $\Phi_{\rm SS}$. For $A>B$, the situation is the opposite. As healing proceeds, $\Phi$ increases and comes closer to $\Phi_{\rm SS}$, which also increases but at a slower pace. Hence, $\Phi$ eventually catches up, achieving a steady state, whereby healing saturates by the convergence to a steady state suitable for the imposed $\tau_{\rm hold}$. Afterward, $\Phi$ remains at that plateau level. This mechanism has been discussed by Ruina (1983) in terms of the stability of the RSF behavior under a constant $\tau$.
@@ A.2 Aging law
which agrees with the experimental relation eq. (67); in Appendix A, variables conditioned by the name of an evolution law denote predictions by the evolution law. We see that the canonical $B\ln t$ healing (eqs. 48 and 49) by converting eq. (A.11) into its $\Phi$-notation form via eq. (11). The predicted healing can be found in subpanels a2 and a3 in Fig. A.1.
@@ A.3 Slip law
that is, when the slip during the hold is relatively small. Note that eq. (A.21) is a subset of the NSC when we assume the canonical $f_{t|{\rm NSC}}$ (eq. 65, §3.3 in the text), as well as $f_{\delta}V=-\Omega\ln\Omega$ of the slip law. Therefore, subsequent predictions in this subsection are all we need to compare against the canonical healing behavior to (in)validate the slip law. Taking the lowest order of $(1-B/A)\delta/D_{\rm c}$, eq. (A.19) reduces to
@@
It is easy to confirm that eq. (A.34) reproduces the level of $\theta$ in the saturation plateau (eq. A.9). Since the slope $B_{\rm heal}$ of the log-t healing predicted by the slip law is $A$ (eq. A.25) instead of the canonical $B$, $t_{*}^{\prime}$ (eq. A.33) is not identical to $t_{*}$ (eq. A.8). Nonetheless, a logical parallel is present between eqs. (A.8) and (A.33); $t_{*}^{\prime}$ in eq. (A.33) is equal to $t_{*}$ in eq. (A.8) after $A$ and $B$ are interchanged, and then eqs. (A.33) and (A.8) are consistent in the lowest-order approximation with respect to $|A-B|/A$.
@@
Equation (A.35) is an order estimate, and thus eq. (A.35) has already been proven for $t\lesssim t_{*}^{\prime}$ through the derivation of eq. (A.25). The following calculation then focuses on the validity of eq. (A.35) for $t\gg t_{*}^{\prime}$. Throughout the remainder of the proof, we suppose $A<B$ and $t\gg t_{*}^{\prime}$, where we derive eq. (A.35) via a perturbation theory by utilizing the fact that healing has proceeded well when $t\gg t_{*}^{\prime}$ (from eqs. A.29 and A.32).
@@ A.4 Composite law
The behavior of the composite law (eq. 74) differs a lot depending on $V_{\rm init}/V_{\rm c}$. The numerical solutions in Fig. A.1(c1-c3) are for $V_{\rm init}/V_{\rm c}\gg 1$, where the predicted healing is much different from the canonical behavior. We start the analysis below with the other, easier case of $V_{\rm init}/V_{\rm c}\ll 1$, where the composite law predicts the canonical healing behavior.
If $V_{\rm init}/V_{\rm c}\ll 1$, this condition will persist forever as $V$ monotonically decreases during constant-$\tau$ healing tests (eq. A.4). Therefore, the composite law is reduced to
@@
Under the presently concerned NSC, $\Omega\ll 1$, so that eq. (A.43) and the aging law are indistinguishable, approximated by $\dot{\theta}\simeq 1$. Then it predicts the canonical healing behavior. Note that, for $A>B$, the healing saturation explained in §A.1 occurs, and the general results discussed there apply. This includes that the strength expected at the saturation plateau coincides with that predicted by the aging law because their predictions of $\Omega_{\rm SS}$ remain constant (eq. 75) throughout the hold period, despite that the constant $\Omega_{\rm SS}$ value expected of eq. (A.43) is $1/W(1)(=1.763...)$, rather than $1$ expected of the aging law. After all, we can conclude
@@
If $V_{\rm init}\gtrsim V_{\rm c}$, the composite law behaves almost like the slip law until the time $t_{\rm sa}$ when $V$ decelerates to $V_{\rm c}$, while it behaves almost like the aging law afterward. We can estimate this switching time $t_{\rm sa}$ by using the slip-law prediction eq. (A.28):
@@
In cases where the healing saturation predicted by the slip law begins earlier than this switch from the slip law to the aging law, the evolution is expected to remain to follow the slip law, and thus healing saturation occurs while still in the slip-law regime; $\Phi$ stops increasing, and in turn, $V$ stops decreasing and stays at a value higher than $V_{\rm c}$. That is, when $A>B\cap t_{\rm sa}\gtrsim t_{*}^{\prime}$, the composite law behaves like the slip law over the entire time range:
@@
In the remaining cases [Fig. A.1(c1-c3) all deal with such cases], the evolution is slip-law-like until $t_{\rm sa}$, after which the evolution is the aging-law-like:
@@
where $\theta_{\rm slip}(t)$ before the saturation time is evaluated with eq. (A.29). For brevity, eq. (A.47) specifies initial ($\theta$, $\Omega$) values in the lower part of the right-hand side (instead of eqs. A.2 and A.6 consistently used in this appendix section, including the left-hand side); the expression for $t\gg t_{\rm sa}$ ($V\ll V_{\rm c}$) considers the fact that $\Omega_{\rm SS}$ changes by a factor of $1/W(1)$ as decreasing $V$ crosses $V_{\rm c}$ from above. Note that because $t_{\rm sa}$ is defined as the time when $V=V_{\rm c}$, eq. (A.3) yields
@@
which is actually satisfied by eqs. (A.29) and (A.45).

### v3

@@ 5.3 Potential detection of $\beta$ from an experiment
As far as $\dot{\theta}=f(V,\theta)$, we have theoretically shown that the aging-slip switch must occur around $\Omega=\beta$ to reproduce the three canonical behaviors confirmed by experiments so far. Such evolution laws involving an aging-slip switch are necessarily accompanied by some side effects violating the canons (§4.1; also, see §4.2.1). Among them, relatively accessible by experiments may be the behavior following a large negative velocity step ($V_{\rm after}/V_{\rm before}\ll\beta$), which is expected to exhibit too rapid strengthening compared to the prediction by the slip law, that is, the VS canon (e.g., Fig. 11a). Although existing VS tests generally obey the slip law as we presumed as the VS canon throughout the present development, a slight deviation has been reported in fact. Bhattacharya et al. (2015) realized that an interesting behavior of the Nagata law (Nagata et al., 2012), which is the aging law plus a stress-dependent term $-cd\tau$ representing an experimentally detected instantaneous effect of the applied shear stress to weaken the interface; that is, the Nagata law changes from the aging law to the slip law for large velocity steps. They further showed that some VS datasets involving multiple steps with various levels of $V_{\rm after}/V_{\rm before}$ (ranging $10^{-2}$ to $10^{2}$) are fit well by the Nagata law with $c=50$, contrasting that the slip law cannot fit the large negative VS. This finding of Bhattacharya et al. (2015) implies that the involvement of the aging-slip switch improves fitting to the experiments.
As Bhattacharya et al. (2015) themselves point out, assuming $c$ as large as 50 is unreasonable as the shear-stress effect of the Nagata law, given that Nagata et al. (2012) obtained $c=$2 by direct measurement of $d\Phi/d\tau$. However, from the viewpoint of the aging-slip switch, their Nagata law fits suggest the aging-slip switch according to $\Omega$ we deduced from the canons. Below, we try fitting the data shown in Bhattacharya et al. (2015) using the proposed modified composite law. Of course, making a formal statistical conclusion on the aging-slip switch according to $\Omega$ requires to demonstrate the increase in likelihoods sufficient to counter the increase in the number of parameters. The slight deviation found in the $V_{\rm after}/V_{\rm before}$ range of the data used in Bhattacharya et al. (2015) is probably not enough to conclude the existence of the additional parameter (excessively large $c$, or reasonably, $\beta$(Bhattacharya et al., 2015). Therefore, a definitive conclusion must be postponed until more suitable data becomes available. Further detailed discussion shall be shown at the end of this subsection, where we also consider Bhattacharya et al. (2022) involving exploration of a further lower range of $\Omega$.
@@
In the above, we have pointed out that the modified-composite law with $\beta=0.01$ explains the data of Bayart et al. (2006) for $V_{\rm after}/V_{\rm before}=0.01$, which Bhattacharya et al. (2015) have found to be basically slip-law-like but somewhat skewed toward aging-law-like behaviors, which are approximated by the Nagata law with excessively large $c$-values. At the same time, we admit that some datasets do not support the aging-slip switch. First, Bhattacharya et al. (2015) note that the above-mentioned better fit of the Nagata law than the slip law, which supports the aging-slip switch, was not evident in one of the three datasets they examined. Second, Bhattacharya et al. (2022) have performed VS tests, planned for capturing the composite-law-like aging-slip switch, strategically including quite low nm/s-order slip rates and small and large steps. Although they have conducted VS tests with $V_{\rm after}/V_{\rm before}$ down to $10^{-3.5}$, they have not reported the aging-slip switch. It implies $\beta$ for their samples is below $10^{-3.5}$ if exits. Still, this observation of Bhattacharya et al. (2022) may originate from the long-term trend (similar to that in Blanpied et al., 1998b) in their data, not fully taken into their analysis. They reported that the observed slip-strengthening behaviors are even slower than the slip law, which would presumably indicate the above-mentioned long-term trend. However, when the long-term trend is removed by a linear fit, we find that the detrended data for $V_{\rm after}/V_{\rm before}=10^{-3.5}$ are faster than the slip law, and the modified composite law with $\beta=0.01$ well fits those detrended data, although the statistical significance of our fitting is yet to be scrutinized. Thus, we deem $10^{-3.5}<\Omega<10^{-2}$ still under debate.
After all, we do not know much about the $\beta$ value, because VS tests with $V_{\rm after}/V_{\rm before}\leq 10^{-2}$ is still scarce. However, we emphasize that the existence of $\beta$ has been deduced from the reconciliation of the three canons under the premise of $f(V,\theta)$, and our SHS canon (significant log-t healing in low-$\tau_{\rm hold}$ SHS) is set at $\Omega$ far-far below steady states ($\Omega_{\rm SS}=1$), that is, $\Omega\to 0$, much lower than $10^{-3.5}$. As the proposers, we reckon that VS tests with further small $V_{\rm after}/V_{\rm before}$ will find the expected switch. Also, note that the value of $\beta$, like the value of $A$ and $B$, can depend on various experimental conditions other than the variables of friction law ($\tau$, $V$, $\Phi$), including rock types, temperature, normal stress, and gouge conditions. Hence, $\beta$ must be considered a fitting parameter, which should be experimentally determined under each condition, as $A$ and $B$ are. For prudence, we still regard the aging-slip switch and its side effects as an unconfirmed theoretical prediction of the present work, and we would like to present the data analysis shown in Fig. 12 as cursory support to our proposal.
@@ 6 Conclusion
In the second step (§4.2.2), we have further shown that the aging-slip switch occurring around $\Omega=\beta$ is an inevitable consequence of the three canons even when evolution laws are sought in a more general scope of $\dot{\theta}=f(V,\theta)=f_{t}+f_{\delta}V$ with the $f=\psi(\Omega)$ ansatz removed (eq. 86).
@@ Acknowledgements.
We would like to thank Dr. Hugo Perfettini, Dr. Paul Antony Selvadurai, Dr. So Ozawa, Dr. Bunichiro Shibazaki, Dr. Yukitoshi Fukahata, Dr. Takeshi Iinuma, and Dr. Takane Hori for valuable discussions and encouragement. We also acknowledge careful reviews and insightful comments from Dr. Allan Rubin, Dr. Shiqing Xu, Dr. Baoning Wu, and an anonymous reviewer. This study was supported in part by JSPS KAKENHI Grant Numbers 22KJ1658, 23K19082, and 21H04507 and by the Ministry of Education, Culture, Sports, Science and Technology (MEXT) of Japan, under its Earthquake and Volcano Hazards Observation and Research Program.
@@ Appendix A Predictions of different evolution laws during quasi-stationary contact
This appendix A presents the behaviors expected of different evolution laws discussed in the present paper (i.e., the aging, slip, composite, and modified composite laws) during constant-$\tau$ SHS tests explained below. In this type of SHS test, following a steady-state sliding at $V=V_{\rm prior}$, shear stress is abruptly reduced by $\Delta\tau$ ($>0$) to a designated level $\tau_{\rm hold}$ ($<\tau_{\rm SS}(V_{\rm prior})$) and held constant for a designated time duration $t_{\rm h}$. Regarding the log-t healing behavior, we focus on constant-$\tau$ SHS tests because the NSC can be ensured throughout the hold period by setting the $\tau_{\rm hold}$ sufficiently low.
For all the four evolution laws considered here (unless $V\lesssim V_{\rm c}$ for the composite law), initial conditions at the beginning of a quasi-stationary hold period (defined as $t=+0$) are given as follows:
$V|_{\rm t=+0}$ $=:V_{\rm init}=V_{\rm prior}e^{-\Delta\tau/A}$ (A.1)
$\theta|_{\rm t=+0}$ $=:\theta_{\rm init}=L/V_{\rm prior}$ (A.2)
Equation (A.1) derives from eq. (1), irrespective of the evolution law here considered. Note that eq. (A.2) assumes each evolution law satisfies $\Omega_{\rm SS}=1$ (eq. 31). This is not always the case, for instance, for the composite law, where $\Omega_{\rm SS}$ is given by $\Omega_{\rm SS}\ln\Omega_{\rm SS}=\exp(-V/V_{\rm c})$. Associated remarks will appear in §A.1 and §A.4. Throughout Appendix A, we also assume $D=L=D_{\rm c}$ and $\theta\gg\theta_{\rm X}$.
@@
where $\Omega_{\rm init}:=\Omega|_{t=+0}$. Equation (A.4) shows that $V$ decreases as the interface strengthens under the constant $\tau_{\rm hold}$ imposed. Equation (A.5) shows that $\Omega$ remains constant if $A=B$, monotonically increases if $A>B$, and monotonically decreases if $A<B$. Using eqs. (A.4) and (A.5), we will analyze the evolution of $\theta$ (and $\Phi$ via eq. 11) for each evolution law in §A.2A.5, after cautioning a problem of constant-$\tau$ SHS tests for interfaces with $A>B$ in §A.1. We focus on behaviors under the NSC by setting $\tau_{\rm hold}$ sufficiently low such that
@@
Figure A.1 summarizes the exact evolutions obtained by numerical integration for all cases of the present concern. We will explain each case through analytical evaluations and discuss how it agrees or disagrees with the canonical behavior of log-t healing (§3.3 and §3.4).
@@ A.1 Healing saturation—a flaw of constant-$\tau_{\rm hold}$ tests for frictional interfaces with $A>B$
In general, $\Phi_{\rm SS}(V)$ increases during a hold period because $V$ decreases as healing proceeds (i.e., $\theta$ increases, eq. A.4). When $A\leq B$, while both $\Phi$ and $\Phi_{\rm SS}$ increases with time, $\Phi$ never catches up with $\Phi_{\rm SS}$; in fact, eq. (A.5) guarantees that $\Phi$ more and more overwhelms $\Phi_{\rm SS}$. For $A>B$, the situation is the opposite. As healing proceeds, $\Phi$ increases and comes closer to $\Phi_{\rm SS}$, which also increases but at a slower pace. Hence, $\Phi$ eventually catches up, achieving a steady state, whereby healing saturates by the convergence to a steady state suitable for the imposed $\tau_{\rm hold}$. Afterward, $\Phi$ remains at that plateau level. This mechanism has been discussed by Ruina (1983) in terms of the stability of the RSF behavior under a constant $\tau$.
@@ A.2 Aging law
which agrees with the empirical relation eq. (67); in Appendix A, variables conditioned by the name of an evolution law denote predictions by the evolution law. We see that the canonical $B\ln t$ healing (eqs. 48 and 49) by converting eq. (A.11) into its $\Phi$-notation form via eq. (11). The predicted healing can be found in subpanels a2 and a3 in Fig. A.1.
@@ A.3 Slip law
that is, when the slip during the hold is relatively small. As shown in §A.3.2, $\Omega\sim 1$ can be realized outside this range so that eq. (A.21) lies in $\Omega\ll 1$, where the canonical $f_{t|{\rm NSC}}$ (eq. 65, §3.3 in the text) and $f_{\delta}V=-\Omega\ln\Omega$ of the slip law contradict each other. Therefore, subsequent predictions in this subsection are all we need to compare against the canonical healing behavior to (in)validate the slip law. Taking the lowest order of $(1-B/A)\delta/D_{\rm c}$, eq. (A.19) reduces to
@@
Equation (A.34) represents $A\ln t$ healing and its saturation at the plateau level of eq. (A.9). Since the slope $B_{\rm heal}$ of the log-t healing predicted by the slip law is $A$ (eq. A.25) instead of the canonical $B$, $t_{*}^{\prime}$ (eq. A.33) is not identical to $t_{*}$ (eq. A.8). Nonetheless, a logical parallel is present between eqs. (A.8) and (A.33); $t_{*}^{\prime}$ in eq. (A.33) is equal to $t_{*}$ in eq. (A.8) after $A$ and $B$ are interchanged, and then eqs. (A.33) and (A.8) are consistent in the lowest-order approximation with respect to $|A-B|/A$.
@@
Equation (A.35) is an order estimate, and thus eq. (A.35) has already been proven for $t\lesssim t_{*}^{\prime}$ through the derivation of eq. (A.25). The following calculation then focuses on the validity of eq. (A.35) for $t\gg t_{*}^{\prime}$. Throughout the remainder of the proof, we suppose $A<B$ and $t\gg t_{*}^{\prime}$, where we derive eq. (A.35) via a perturbation theory by utilizing the fact that healing has proceeded well when $t\gg t_{*}^{\prime}$ (eqs. A.29 and A.32).
@@ A.4 Composite law
The behavior of the composite law (eq. 74) differs a lot depending on $V_{\rm init}/V_{\rm c}$. The numerical solutions in Fig. A.1 (c1-c3) are for $V_{\rm init}/V_{\rm c}\gg 1$, where the predicted healing significantly differs from the canonical log-t healing. We start the analysis below with the other, more straightforward case of $V_{\rm init}/V_{\rm c}\ll 1$, where the composite law reproduces the SHS||||NSC canon.
If $V_{\rm init}/V_{\rm c}\ll 1$, $V/V_{\rm c}\ll 1$ holds forever as $V$ monotonically decreases during constant-$\tau$ healing tests (eq. A.4). Therefore, the composite law is reduced to
@@
Under the presently concerned NSC, $\Omega\ll 1$ so that eq. (A.43) is approximated by $\dot{\theta}\simeq 1$, reproducing the canonical log-t healing. Note that, for $A>B$, the healing saturation explained in §A.1 occurs given the constancy of $\Omega_{\rm SS}$ (eq. 75) throughout the hold period; the composite-law prediction of the plateaued strength value may be different from that of the aging law because the $\Omega_{\rm SS}$ value expected of eq. (A.43) is $1/W(1)(=1.763...)$, rather than unity expected of the aging law, while the composite law predicts $\Omega_{\rm init}=e^{-\Delta\tau/A}$ as in the aging law for ordinary cases of $V_{\rm prior}\gg V_{\rm c}$. After all, we can conclude
@@
If $V_{\rm init}\gtrsim V_{\rm c}$, the composite law behaves almost like the slip law until the time $t_{\rm sa}$ when $V$ decelerates to $V_{\rm c}$, while it behaves almost like the aging law afterward. We can estimate this switching time $t_{\rm sa}$ by using the slip-law prediction of $V$ (eq. A.28):
@@
In cases where the healing saturation predicted by the slip law begins earlier than this switch from the slip law to the aging law, the evolution remains to follow the slip law, and thus healing saturation occurs while still in the slip-law regime; $\Phi$ stops increasing, and in turn, $V$ stops decreasing and stays at a value higher than $V_{\rm c}$. That is, when $A>B\cap t_{\rm sa}\gtrsim t_{*}^{\prime}$, the composite law behaves like the slip law over the entire time range:
@@
In the remaining cases [Fig. A.1 (c1–c3) all deal with such cases], the evolution is slip-law-like until $t_{\rm sa}$, after which the evolution is the aging-law-like:
@@
where $\theta_{\rm slip}(t)$ before the crossover is evaluated with eq. (A.29). For brevity, eq. (A.47) specifies initial ($\theta$, $\Omega$) values in the lower part of the right-hand side (instead of eqs. A.2 and A.6 consistently used in this appendix section including the left-hand side of eq. A.47); eq. (A.47) for $t\gg t_{\rm sa}$ ($V\ll V_{\rm c}$) accounts for the fact that $\Omega_{\rm SS}$ changes by a factor of $1/W(1)$ as the decreasing $V$ crosses $V_{\rm c}$ from above. Note that because $t_{\rm sa}$ is defined as the time when $V=V_{\rm c}$, eq. (A.3) yields
@@
which is certainly satisfied by eqs. (A.29) and (A.45).